The manager of a garden shop mixes grass seed that is 30% rye grass with 56 lb of grass seed that is 75% rye grass to make a mixture that is 65% rye grass. How much of the 30% rye grass is used?
=______LB

Answer by Theo(13342)

let x = amount of grass that is 30% rye.
56 = amount of grass that is 75% rye.
x + 56 = amount of grass that is 65% rye.
you get:
.3 * x + .75 * 56 = .65 * (x + 56)
simplify to get:
.3 * x + 42 = .65 * x + 36.4
subtract .3 * x from both sides of the equation and subtract 36.4 from both sides of the equation to get:
42 - 36.4 = .65 * x - .3 * x
simplify to get:
5.6 = .35 * x
solve for x to get:
x = 5.6 / .35 = 16
your solution is that 16 pounds of the 30% rye grass is used.